# The causal structure of galactic astrophysics

### v1

@@ Abstract
keywords:
methods: data analysis – methods: statistical – galaxies: formation – galaxies: fundamental parameters – galaxies: statistics
pubyear: 2025pagerange: The causal structure of galactic astrophysicsLABEL:LastPage
@@ 1 Introduction
Understanding the physical processes that shape galaxies is a central goal of astrophysics. Empirical progress has traditionally relied on identifying correlations between observed properties, which can then be interpreted in light of theoretical models for galaxy formation and used to constrain them. The advent of large surveys and powerful machine learning techniques has greatly expanded our ability to find such statistical associations, uncovering intricate patterns across high-dimensional parameter spaces. However, correlation alone is insufficient for determining causal influences among variables: which properties are actually responsible for determining others, in what direction this influence goes, and whether there exist confounding variables that are not included in the dataset but influence those that are. Achieving this requires causal discovery, a methodology widely applied in fields such as genomics, epidemiology and economics, but that has had extremely limited exposure in the physical sciences.
@@
A possibly complete list of previous applications of causal discovery to astrophysics follows. Mucesh et al. (2024) estimate a causal model of galaxy formation from semi-analytic models and hydrodynamical simulations, and compare it to non-causal models. Pasquato et al. (2023) apply the Peter–Clark and Fast Causal Inference algorithms to 83 galaxies in an attempt to constrain evolution mechanisms for their central supermassive black holes. Jin et al. (2024); Jin et al. (2025b) addresses the same question with 101 galaxies, using a Bayesian method for estimating the probabilities of all possible causal structures. Finally, Jin et al. (2025a) apply a linear causal discovery model to 100 simulated galaxies to constrain chemodynamical pathways relevant for the Milky Way. These studies involve very small numbers of galaxies far from representative of the population at large, and have been restricted to very specific variable sets and scientific questions.
After describing causal discovery and identifying a method suitable for astrophysical problems, we illustrate the technique with galaxy data. Specifically we take seven variables describing the first-order properties (brightness, mass, size, morphology, star formation rate) of $\sim 5\times 10^{5}$ low-redshift galaxies from the Nasa Sloan Atlas (NSA). We assess reliability of the causal discovery and calibrate hyperparameters of our algorithm on mock data, then infer the causal links present in the NSA including the presence of confounding latent variables. We show explicitly how this helps to pinpoint the physical mechanisms governing the data, which are crucially directional, in addition simply to inducing correlations.
@@ 2 Observational data
We base our analysis on the NSA v1.0.1 (Blanton et al., 2011),111www.sdss4.org/dr17/manga/manga-target-selection/nsa/ a value-added catalogue of nearby galaxies that includes inferred quantities (such as stellar mass and star formation rate) in addition to raw observables. It is based primarily on Sloan Digital Sky Survey (SDSS) imaging but employs reprocessed reductions with improved sky subtraction and photometry tailored for extended sources. The catalogue includes galaxies with redshifts $z\lesssim 0.15$, and provides homogenised multi-band photometry and spectroscopic redshifts, including a cross-match with Galaxy Evolution Explorer (GALEX) data to fill in the ultra-violet part of galaxies’ spectral energy distributions. The NSA is a widely-studied standard for local galaxies (e.g. Reines et al. 2013; Wheeler et al. 2014; Ma et al. 2014; Bundy et al. 2015; Latimer et al. 2021).
@@
  • zdist: estimated cosmological redshift correcting the observed redshift with a peculiar velocity estimate. This corresponds approximately to distance in Mpc/$h$.
@@
  • elpetro_b300: current star formation rate (SFR) divided by the average over the past 300 Myrs.
@@
These are designed to capture the most important information about galaxy structure, including mass, luminosity, size, structure and redshift. The quantities designated ‘elpetro’ derive from elliptical Petrosian flux fits, which are deemed the most reliable in the catalogue. elpetro_mass and elpetro_b300 have been $K$-corrected to rest-frame magnitudes using the kcorrect code (Blanton & Roweis, 2007). Absolute magnitudes are given with $H_{0}=100h$ km/s/Mpc so should be read as $M-5\log{h}$. All logs have base 10.
@@ 3.1 Causal Discovery
Most statistical analyses in astrophysics (whether or not through the lens of machine learning) are designed to measure correlations: how strongly two quantities co-vary and the properties of their relationship. Correlation, however, is agnostic about direction and mechanism, which are the predictions of galaxy formation theories and hence the most useful features for testing them. These predictions can be projected onto the space of correlations, but information and hence constraining power is lost in doing so. Causal discovery methods seek retain this information, going beyond correlation by inferring the causal structure that generates the observed data. This provides added value to the results that complements or is overlaid upon the traditional results of astrostatistical methodology. For thorough reviews of causal discovery see Spirtes et al. (1993); Mooij et al. (2014); Eberhardt (2017); Glymour et al. (2019).
@@
Causal discovery utilises conditional correlation strengths to uncover the directions of influence among variables. To see how this is possible, suppose we measure three galaxy properties: stellar mass ($M$), gas mass ($G$), and star-formation rate ($S$). Observationally we find that all three are correlated: larger values of one are associated with larger values of any other. From this alone, one could imagine a mass-driven scenario (more massive galaxies accrete more gas, $M\rightarrow G$, which in turn fuels star formation, $G\rightarrow S$), a feedback-driven scenario (high $S$ regulates gas supply, $S\rightarrow G$, while simultaneously building stellar mass, $S\rightarrow M$) or a common-cause scenario (the environment controls both mass growth and gas supply, indirectly correlating all three). However, if $M$ and $S$ remain correlated even after conditioning on $G$, there must exist a direct causal link between them ($M\rightarrow S$); conversely if conditioning on $G$ removes the MSM-S correlation we must instead have $M\rightarrow G\rightarrow S$. Thus the existence or absence of the conditional correlation breaks the degeneracy between the physical scenarios.
@@
By assessing all such conditional correlations (including a multi-dimensional conditioning set) one can determine the Markov equivalence class to which the data belongs. Each such class contains the set of causal structures with the same statistical dependencies, and which therefore cannot be distinguished without experimental intervention, impossible in astrophysics. (Such statistical independencies can also be thought of as implying factorisability of the joint (probability) distribution describing the variables: for example $A\rightarrow B\rightarrow C$ implies $P(A,B,C)=P(A)P(B|A)P(C|B)$.) This leads to the classic constraint-based causal discovery method, the Peter–Clark algorithm, which eliminates direct (i.e. causal) correlations with conditional independence tests then applies orientation rules to fill in directions where possible. [^2] As an example of such a rule, suppose that one has identified the direct links ABA-B, ACA-C, BDB-D, CDC-D. If $B$ and $C$ are disconnected when conditioning on $A$ but not when conditioning on $D$, it must be that neither $B$ nor $C$ are caused by $D$. Hence the BDB-D and CDC-D links must be $B\rightarrow D$ and $C\rightarrow D$. Alternative score-based methods such Greedy Equivalent Search assign likelihood scores to candidate solutions based on the correlation strengths and search for the highest-scoring structures, while functional causal models such as additive noise models instead exploit asymmetries in functional relationships to determine causal direction.
@@
Causal structure is visualised in a causal graph—a directed network of relationships among the variables. The true, generating causal structure is described by a directed acyclic graph (DAG), in which all edges (i.e. causal links between variables, represented by lines) are directed. The corresponding Markov equivalence class is shown by a completed partially directed acyclic graph (CPDAG), which contains edges for which directionality cannot be established based solely on the conditional independencies. A partial ancestral graph (PAG) additionally models the potential effect of latent confounding variables. These can create dependencies between observables that cannot be resolved by any orientation of arrows among the observed variables alone: every candidate orientation produces an inconsistency with other independencies. This replaces the DAG with a maximal ancestral graph (MAG), indicating the possible influence of hidden causes with a circle endpoint. A PAG then represents the Markov equivalence class of the true MAG, and is the goal of algorithms that do not assume causal sufficiency.
@@
Several assumptions are required for causal discovery to be possible. The most common are the Markov condition (separated variables in the causal graph are statistically independent), faithfulness (no accidental statistical independences) and acyclicity (nothing can be indirectly its own cause). In addition, the conditional correlations must be adequately captured by the statistical test applied (which come with their own assumptions) and the threshold $p$-value chosen to weed out insignificant correlations.
@@ 3.2 The FCIT algorithm applied to galaxies
For application to the NSA, we have the requirements that a method is 1) accurate in the presence of confounding latent variables, since it is highly unlikely that all relevant information is contained in the dataset, 2) able to accommodate highly non-linear correlations (see Fig. 1), and 3) efficient enough to analyse $\mathcal{O}(10^{5})$ objects in reasonable time.
To achieve this we adopt the newly-developed method Fast Causal Inference with Targeted Testing (FCIT; Ramsey et al 2025, in prep) as implemented in py-tetrad (Ramsey & Andrews, 2023). [^3] https://www.cmu.edu/dietrich/philosophy/tetrad/use-tetrad/tetrad-python.html, [^4] https://github.com/cmu-phil/py-tetrad This is a hybrid constraint-and-score-based algorithm which starts with a score-based estimate of the causal graph and then identifies a minimal set of conditional independence tests required to remove a causal link between two variables, given the provisional PAG. This reduces the number of tests required and the statistical noise they introduce. It also incorporates discriminating path checks during edge removal, ensuring that edges are properly oriented before deciding on conditional independence. The resulting graphs are edge-minimal, correctly oriented and exhibit high accuracy on causal discovery benchmarks, with a runtime of only $\sim$1 minute for $5\times 10^{5}$ datapoints and 7 features.
@@
where $\mathcal{L}$ is the likelihood of a target variable when predicted according to a hypothetical conditional correlation structure, $N$ is the sample size used for the local regression, and $k$ is the number of free regression coefficients in the basis‑function expansion for the child given its current parent set.
The two main hyperparameters affecting dataset-specific performance are penalty_discount and truncation_limit. The former controls how strongly graph complexity is penalised (Eq. 1). A higher value favours simpler graphs by removing more noise-dominated edges, at the expense of the quality of the conditional fit. The latter controls the complexity of the local regression model through the number of polynomial basis terms included. Larger values allow more expressive models at the cost of runtime and reduced BIC.
@@ 3.3 Mock data generation
To optimise these hyperparameters for our astrophysical application we create mock datasets with similar characteristics to the real data but with known causal structure. This will also enable the reliability of the method to be quantified.
This is achieved with the Causal Perceptron Network (CPN; Ramsey et al 2025, in prep), a code for generating synthetic datasets from arbitrary nonlinear models. The user specifies a DAG that encodes the desired causal structure, along with a noise distribution. Each variable is then expressed as a nonlinear function of its causes plus an independent noise term. Rather than choosing simple algebraic forms for these functions, CPN uses randomly configured multilayer perceptrons. Each dataset is made by recursively sampling noise and propagating values forward through the causal graph, producing independent and identically distributed samples. This lets CPN produce highly flexible data approximating a broad class of nonlinear functions.
@@ 4.1 Mock data tests
We find that truncation_limit =14=14 is ideal for this data: it is considerably larger than the default value of 3, allowing the highly nonlinear relations between variables to be captured, but still larger values tend to decrease the BIC due to the additional model complexity. The results are largely insensitive to this. Fixing this we then scan through penalty_discount, calculating in each case the average precision, recall and F1 score across 200 mock datasets differing only in their number of edges and the random number generation.
The result is shown in Fig. 2. As penalty_discount increases, the scoring function penalises model complexity more strongly, leading to sparser graphs. This reduces false positives and thus tends to increase precision, but it also causes some true edges to be missed, lowering recall. The F1 score exhibits a peak at penalty_discount $\sim 40-50$ at a value $\sim 0.9$, roughly indicating a 90 per cent success rate on each dataset. We adopt a value of 50 for the real data, but again explore reasonable variations without finding important differences. We also confirm that the other hyperparameters in the FCIT algorithm and its testing and scoring methods have little impact on the results.

### v2

@@ 1 Introduction
Understanding the physical processes that shape galaxies is a central goal of astrophysics. Empirical progress has traditionally relied on identifying correlations between observed properties, which can then be interpreted in light of theoretical models for galaxy formation and used to constrain them. The advent of large surveys and powerful machine learning techniques has greatly expanded our ability to find such statistical associations, uncovering intricate patterns across high-dimensional parameter spaces. However, correlation alone cannot determine causal influences among variables: which properties are actually responsible for determining others, in what direction this influence goes, and whether there exist confounding variables that are not included in the dataset but influence those that are. Achieving this requires causal discovery, a methodology widely applied in fields such as genomics, epidemiology and economics, but that has had extremely limited exposure in the physical sciences.
@@
A possibly complete list of previous applications of causal discovery to astrophysics follows. Mucesh et al. (2024) estimate a causal model of galaxy formation from semi-analytic models and hydrodynamical simulations, and compare it to non-causal models. Pasquato et al. (2023) apply the Peter–Clark and Fast Causal Inference algorithms to 83 galaxies in an attempt to constrain evolution mechanisms for their central supermassive black holes. Jin et al. (2024); Jin et al. (2025b) addresses the same question with 101 galaxies, using a Bayesian method for estimating the probabilities of all possible causal structures. Davis et al. (2025) use Fast Causal Inference to help constrain the formation histories of trans-Neptunian objects. Finally, Jin et al. (2025a) apply a linear causal discovery model to 100 simulated galaxies to constrain chemodynamical pathways relevant for the Milky Way. These studies involve samples too small to be representative of the underlying populations, and have been restricted to very specific variable sets and scientific questions.
After describing causal discovery and identifying a method suitable for astrophysical problems involving hundreds of thousands of objects, we illustrate the technique with galaxy data. Specifically we take seven variables describing the first-order properties (brightness, mass, size, morphology, star formation rate) of $\sim 5\times 10^{5}$ low-redshift galaxies from the Nasa Sloan Atlas (NSA). We assess reliability of the causal discovery and calibrate hyperparameters of our algorithm on mock data, then infer the causal links present in the NSA including the presence of confounding variables not in the data subset. We show explicitly how this can be used to refine our understanding of galaxy evolution and test proposed physical mechanisms, which crucially (as is typical in physics) come with a direction of operation in addition to simply inducing correlation among galaxy properties.
@@ 2 Observational data
We base our analysis on the NSA v1.0.1 (Blanton et al., 2011),111www.sdss4.org/dr17/manga/manga-target-selection/nsa/ a value-added catalogue of nearby galaxies that includes inferred quantities (such as stellar mass and star formation rate) in addition to raw observables. It is based primarily on Sloan Digital Sky Survey (SDSS) imaging but employs reprocessed reductions with improved sky subtraction and photometry tailored for extended low-redshift sources. The catalogue includes galaxies with redshifts $z\lesssim 0.15$, and provides homogenised multi-band photometry and spectroscopic redshifts, including a cross-match with Galaxy Evolution Explorer (GALEX) data to fill in the ultra-violet part of galaxies’ spectral energy distributions. The NSA is a widely-studied standard for the local galaxy population (e.g. Reines et al. 2013; Ma et al. 2014; Wheeler et al. 2014; Bundy et al. 2015; Latimer et al. 2021).
@@
  • zdist: estimated cosmological redshift correcting the observed redshift with a peculiar velocity estimate from Willick et al. (1997). This corresponds approximately to distance in Mpc/$h$.
@@
  • elpetro_b300: current star formation rate (SFR) divided by the average over the past 300 Myr.
@@
These are designed to capture the most important information about galaxy structure, including mass, luminosity, size, structure and redshift. Note that they are by no means unique: different choices would reflect different user priorities and beliefs about the most causally relevant variables. The quantities designated ‘elpetro’ derive from elliptical Petrosian flux fits, which are deemed the most reliable in the catalogue. elpetro_mass and elpetro_b300 have been $K$-corrected to rest-frame magnitudes using the kcorrect code (Blanton & Roweis, 2007). Absolute magnitudes are given with $H_{0}=100h$ km/s/Mpc so should be read as $M-5\log{h}$. All logarithms are base 10.
@@ 3.1 Causal Discovery
Most statistical analyses in astrophysics (whether or not through the lens of machine learning) are designed to measure correlations: how strongly two quantities co-vary and the properties of their relationship. Correlation, however, is agnostic about direction and mechanism, which are the predictions of galaxy formation theories and hence the most useful features for testing them. Projecting these predictions onto the space of correlations loses information and hence constraining power. Causal discovery methods seek to retain this information, going beyond correlation by inferring the causal structure that generates the observed data. This provides added value to the results that complements or is overlaid upon the traditional results of astrostatistical methodology. It is this information that connects directly to theories and models of galaxy formation and evolution. For thorough reviews of causal discovery see Spirtes et al. (1993); Mooij et al. (2014); Eberhardt (2017); Glymour et al. (2019).
@@
Causal discovery utilises conditional correlation strengths to uncover the directions of influence among variables. To see how this works, suppose we measure three galaxy properties: stellar mass ($M$), gas mass ($G$), and star-formation rate ($S$). We find all three are correlated: larger values of one are associated with larger values of any other. From this alone, one could imagine a mass-driven scenario (more massive galaxies accrete more gas, $M\rightarrow G$, which in turn fuels star formation, $G\rightarrow S$), a feedback-driven scenario (high $S$ regulates gas supply, $S\rightarrow G$, while simultaneously building stellar mass, $S\rightarrow M$) or a common-cause scenario (the environment controls both mass growth and gas supply, indirectly correlating all three). However, if $M$ and $S$ remain correlated even after conditioning on $G$, there must exist a direct causal link between them ($M\rightarrow S$), while if conditioning on $G$ removes the MSM-S correlation we must instead have $M\rightarrow G\rightarrow S$. Thus the existence or absence of the conditional correlation breaks the degeneracy between the physical scenarios.
@@
By assessing all such conditional correlations (including a multi-dimensional conditioning set) one can determine the Markov equivalence class to which the data belongs. Each class comprises causal structures sharing the same statistical dependencies which therefore cannot be distinguished without experimental intervention, impossible in astrophysics. (Such statistical independencies can also be thought of as implying factorisability of the joint (probability) distribution describing the variables: e.g. $A\rightarrow B\rightarrow C$ implies $P(A,B,C)=P(A)P(B|A)P(C|B)$.) This leads to the classic constraint-based causal discovery method, the Peter–Clark algorithm, which eliminates direct (i.e. causal) correlations with conditional independence tests then applies orientation rules to fill in directions where possible.  [^2] As an example of such a rule, suppose that one has identified the direct links ABA-B, ACA-C, BDB-D, CDC-D. If $B$ and $C$ are disconnected when conditioning on $A$ but not when conditioning on $D$, it must be that neither $B$ nor $C$ are caused by $D$. Hence the BDB-D and CDC-D links must be $B\rightarrow D$ and $C\rightarrow D$. There are two more rules involving collisions that can be used to determine directions. Alternative score-based methods such Greedy Equivalent Search assign likelihoods to candidate solutions based on the correlation strengths and search for the highest-scoring structures, while functional causal models such as additive noise models instead exploit asymmetries in functional relationships to determine causal direction.
@@
Causal structure is visualised in a causal graph—a directed network of relationships among the variables. The true, generating causal structure is described by a directed acyclic graph (DAG), in which all edges (i.e. causal links between variables, represented by lines) are directed. The corresponding Markov equivalence class is shown by a completed partially directed acyclic graph (CPDAG), which contains edges for which directionality cannot be established based solely on the conditional independencies. A partial ancestral graph (PAG) additionally models the potential effect of latent confounding variables, overcoming the assumption that all causes are contained within the dataset. These can create dependencies between observables that cannot be resolved by any orientation of arrows among the observed variables alone: every candidate orientation produces an inconsistency with other independencies. This replaces the DAG with a maximal ancestral graph (MAG). A PAG then represents the Markov equivalence class of the true MAG, indicating the possible influence of hidden causes with a circle endpoint. This is the goal of causal discovery.
@@
Several assumptions are required for causal discovery to be possible. The most common are the Markov condition (separated variables in the causal graph are statistically independent), faithfulness (no accidental statistical independences) and acyclicity (nothing can be indirectly its own cause). In addition, the conditional correlations must be adequately captured by the statistical test applied (which come with their own assumptions) and the threshold $p$-value chosen to identify insignificant correlations. Some methods further assume causal sufficiency, namely that all causally relevant variables are included in the dataset. Since causal discovery cannot guarantee the “true” causal graph, but rather identifies the set of structures compatible with the observed data and assumptions, it is best used to weed out statistically implausible causal relationships and hence generate testable causal hypotheses to be compared with theoretical expectations or followed up with targeted observations.
@@ 3.2 The FCIT algorithm applied to galaxies
For application to the NSA, we have the requirements that a method is 1) accurate in the presence of confounding latent variables, since it is highly unlikely that all relevant information is contained in the dataset, 2) able to accommodate highly non-linear functional relationships (see Fig. 1), and 3) efficient enough to analyse $N\approx\mathcal{O}(10^{5})$ objects in reasonable time. This is impossible for traditional algorithms, which scale as $N^{3}$ or worse when accounting for confounders and nonlinearity and become prohibitively memory intensive with $\gtrsim 500$ samples (Strobl et al., 2017; Raghu et al., 2018).
To achieve this we adopt the newly-developed method Fast Causal Inference with Targeted Testing (FCIT; Ramsey et al. 2025a) as implemented in py-tetrad (Ramsey & Andrews, 2023). [^3] https://www.cmu.edu/dietrich/philosophy/tetrad/use-tetrad/tetrad-python.html, [^4] https://github.com/cmu-phil/py-tetrad This is a hybrid constraint-and-score-based algorithm which starts with a score-based estimate of the causal graph, possibly with additional adjacencies that the score-based method cannot remove directly, and then applies a recursive procedure to identify minimal sets of conditional independence tests for removing these extra causal links between variables that should be non-adjacent in the correct PAG. This reduces the number of tests required and the statistical noise they introduce. It also incorporates discriminating path checks during edge removal, ensuring that edges are properly oriented before deciding on conditional independence. Additional bookkeeping in the FCIT procedure guarantees that interim and final graphs are guaranteed to be well-formed PAGs when applied to data samples, a departure from existing latent variable procedures that yield PAGs theoretically. The resulting graphs are edge-minimal, correctly oriented and exhibit high accuracy on causal discovery benchmarks, with an unprecedented runtime of only $\sim$1 minute for $5\times 10^{5}$ datapoints and 7 features.
@@
where $\mathcal{L}$ is the likelihood of a target variable according to a hypothetical conditional correlation structure, $N$ is the sample size used for the local regression, and $k$ is the number of free regression coefficients in the basis‑function expansion for the child given its current parent set (Ramsey et al., 2025b). This makes FCIT a hybrid algorithm: the constraint phase uses the basis‑function likelihood ratio test to prune edges, while the score phase (and the refinement and orientation phases) uses the basis‑function BIC. For the $p$-value threshold we use 0.01 in all cases, which we find not to affect the results appreciably.
Two main hyperparameters affect performance in a dataset-specific way: penalty_discount and truncation_limit. The former controls how strongly graph complexity is penalised (Eq. 1). A higher value favours simpler graphs by removing more noise-dominated edges, at the expense of the quality of the conditional fit as described by $\mathcal{L}$. The latter controls the complexity of the local regression model through the number of polynomial basis terms included. Larger values allow more expressive models at the cost of runtime and enhanced BIC complexity penalty.
@@ 3.3 Mock data generation
To optimise these hyperparameters for our astrophysical application we create mock datasets with similar characteristics to the real data but with known causal structure. This will also enable the reliability of the method to be quantified, giving an indication of the systematic uncertainty possible in the results on the real data.
This is achieved with the Causal Perceptron Network (CPN; Ramsey et al. 2025b), a simulation framework for generating synthetic datasets from arbitrary nonlinear models. The user specifies a DAG that encodes the desired causal structure, along with a noise distribution. Each variable is then expressed as a nonlinear function of its causes plus an independent noise term. Rather than choosing simple algebraic forms for these functions CPN uses randomly configured multilayer perceptrons, ensuring that the resulting data exhibits realistic nonlinear dependencies while remaining stable even with a large number of variables. Each dataset is made by recursively sampling noise and propagating values forward through the causal graph, producing independent and identically distributed samples. This lets CPN produce highly flexible data approximating a broad class of nonlinear functions.
@@ 4.1 Mock data tests
We find that truncation_limit =14=14 is ideal for this data: it is considerably larger than the default value of 3, allowing the highly nonlinear relations between variables to be captured in the FCIT scoring and conditional independence tests, but still larger values tend to decrease the BIC due to the additional model complexity. The results are in any case largely insensitive to this. Fixing this we then scan through penalty_discount, calculating in each case the average precision, recall and F1 score across 200 mock datasets differing only in their number of edges and the random number generation.
The result is shown in Fig. 2. As penalty_discount increases the score penalises model complexity more strongly, leading to sparser graphs. This reduces false positives and thus tends to increase precision, but also causes some true edges to be missed, lowering recall. The F1 score peaks at penalty_discount $\sim 40-50$ at a value $\sim 0.9$, roughly indicating a 90 per cent success rate on each dataset. We adopt a value of 50 for the real data, but again explore reasonable variations without finding important differences. We also confirm that the other hyperparameters in the FCIT algorithm and its testing and scoring methods impact the results little.